2 NO(g) O2 (g) 2NO2 (g) Express your answer using
Answer #1

For:

4N2O --> 3N2 + 2NO2

apply:

multiply equation (3) by 2, since we need N2O2 with 4 coefficient

4N2O = 4N2 + 2O2 dG = (2)(-207.4) = -414.8 kJ/mol

2NO + O2 = 2NO2 dG = -71.2 kJ/mol

Add both equations

4N2O +2NO + O2 = 4N2 + 2O2 +2NO2 dG = -414.8 -71.2 = -486kJ/mol

4N2O +2NO = 4N2 + O2 +2NO2    dG = -414.8 -71.2 = -486kJ/mol

we need to get rid or NO, so add equaiton 2

4N2O +2NO = 4N2 + O2 +2NO2    dG = -414.8 -71.2 = -486kJ/mol

N2 + O2 = 2NO dG = 175.2

add all

N2 + O2 + 4N2O +2NO = 4N2 + O2 +2NO2+ 2NO     dG =-486kJ/mol +175.2

4N2O = 3N2 + +2NO2 dG =-486 +175.2 = -310.8 kJ/mol

so..

dG = -310.8 kJ/mol

B)

2NO + O2 = 2NO2 dG = -71.2 kJ

so..

dG° = -RT*ln(K)

K = exp(-dG°/(RT))

K = exp(71200/(8.314*298)) = 3.024*10^12